OpenStudy (anonymous):

Which of the following lines has a slope of -1/2? x + 2y = 0 x - 2y = 0 -x + 2y = 0

OpenStudy (anonymous):

x + 2y = 0

OpenStudy (mathmale):

I'd suggest you solve each of the three given equations for y (it won't take long). Then the coefficient of x is the slope you're looking for.

OpenStudy (anonymous):

steps 1. x+2y=0 2. 2y=-x+0 3. y=-1/2x